# ReNiL: Relative Neural Inertial Locator with Any-Scale Bayesian Inference

### v2

@@ V-E Application Performance
We implemented the construction of behavioral semantic landmarks based on the approach introduced by NIU [42], with two key modifications: instead of using the step-by-step window integration method for displacement estimation, we employed our IPDPs estimation framework. We incorporated the motion-aware orientation filter to ensure that the constructed map’s coordinate system aligns with the ENU reference frame.
@@
VI Conclusion
This paper introduces ReNiL, a Bayesian inertial-localization framework that shifts the focus from dense trajectory tracking to accurate position and uncertainty estimation at IPDPs. ReNiL combines the motion-aware orientation filter with ASLE, enabling robust displacement and homogeneous uncertainty regression for IMU sequences of any scale. ReNiL also provides a general Bayesian process that accommodates standalone inertial use or loose coupling with external sensors. Future work will explore (i) cross-device and cross-activity generalization without domain-specific fine-tuning, (ii) real-time deployment on ultra-low-power edge processors, and (iii) tighter fusion with GNSS or other signals to further improve availability in challenging environments. We believe ReNiL and the IPDPs perspective establish a solid foundation for next-generation pedestrian localization services that are accurate, efficient, and uncertainty-aware.

### v3

@@ 5.4 Application Performance in Real World
We implemented the construction of behavioral semantic landmarks based on the approach introduced by NIU [niu2020atlas, 10.1145/3749455], with two key modifications: instead of using the step-by-step window integration method for displacement estimation, we employed our IPDPs estimation framework. We incorporated the motion-aware orientation filter to ensure that the constructed map’s coordinate system aligns with the ENU reference frame.
@@
6 Conclusion
This paper introduces ReNiL, an event-driven Bayesian inertial-localization framework that shifts the focus from dense trajectory tracking to accurate position and uncertainty estimation at IPDPs. ReNiL combines the motion-aware orientation filter with ASLE, enabling robust displacement and homogeneous uncertainty regression for IMU sequences of any scale. ReNiL also provides a general Bayesian process that accommodates standalone inertial use or loose coupling with external sensors. Future work will explore (i) cross-device and cross-activity generalization without domain-specific fine-tuning, (ii) real-time deployment on ultra-low-power edge processors, and (iii) tighter fusion with GNSS or other signals to further improve availability in challenging environments. We believe ReNiL and the IPDPs perspective establish a solid foundation for next-generation pedestrian localization services that are accurate, efficient, and uncertainty-aware.